OpenStudy (anonymous):

find the derivatives : d/dx[(x^2+1)sin√x]

OpenStudy (lgbasallote):

\[(x^2 + 1) \sin \sqrt x\] use product rule \[(x^2 + 1) \frac{d}{dx} (\sin \sqrt x) + \sin \sqrt x \frac{d}{dx} (x^2 + 1)\]

OpenStudy (lgbasallote):

so where are you having trouble?

OpenStudy (anonymous):

i dont even know where to start :/

OpenStudy (cherylim23):

You also need the chain rule for d/dx sin( sqrt (x))

OpenStudy (anonymous):

so how do i do that?

OpenStudy (lgbasallote):

first of all do you know product rule?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so do you agree with my set up above?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so your only problem is \[\frac{d}{dx} \sin \sqrt x\] and \[\frac{d}{dx} (x^2 + 1)\]

OpenStudy (lgbasallote):

so what's \[\frac{d}{dx} \sin \sqrt x\]

OpenStudy (anonymous):

i got the same thing that she did

OpenStudy (anonymous):

and now it is gone

OpenStudy (anonymous):

but isnt it cos sqrtx

OpenStudy (lgbasallote):

something is missing

OpenStudy (lgbasallote):

you need to get the derivative of sqrt x furthermore

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

what is the derivative of sqrt x?

OpenStudy (anonymous):

x^-1/2

OpenStudy (anonymous):

is that right?

OpenStudy (lgbasallote):

no..what did you do?

OpenStudy (anonymous):

oh i for got to multiply the front

OpenStudy (anonymous):

1/2 x ^-1/2

OpenStudy (lgbasallote):

better

OpenStudy (anonymous):

ok so how does that go into the problem?

OpenStudy (lgbasallote):

so the \[\frac{d}{dx}\sin \sqrt x \implies \cos \sqrt x (\frac{1}{2\sqrt x}\]

OpenStudy (lgbasallote):

or in nicer words \[\frac{\cos \sqrt x}{2\sqrt x}\]

OpenStudy (anonymous):

ok so where are we in the problem so far?

OpenStudy (anonymous):

like from the beginning

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

now you have to find \[\frac{d}{dx} (x^2 + 1)\]

OpenStudy (anonymous):

ok so thats not my answer is it? sorry im really confused

OpenStudy (lgbasallote):

it's not

OpenStudy (lgbasallote):

remember we have \[(x^2 + 1) (\frac{d}{dx}[ \sin \sqrt x) + \sin \sqrt x (\frac{d}{dx} (x^2 + 1))\]

OpenStudy (lgbasallote):

we already hae d/dx sin sqrt x

OpenStudy (anonymous):

ok so where do we go from there.. i think it might be easier if you would write the whole process out and then explain it as a look at it if you could do that? sorry im just very visual learner

OpenStudy (lgbasallote):

from where?

OpenStudy (anonymous):

from the last thing that you wrote the long equation

OpenStudy (lgbasallote):

that was just the thingy a while ago \[\frac{d}{dx} [(x^2+1)(\sin \sqrt x)] = (x^2 + 1)[\frac{d}{dx} (\sin \sqrt x)] + (\sin \sqrt x)[\frac{d}{dx} (x^2 +1 )]\] that's just the product rule

OpenStudy (lgbasallote):

the first step to solving this is \[\frac{d}{dx} (\sin \sqrt x)\] we already did this correct?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

btw...that's the first step because that's the thing we dont know

OpenStudy (lgbasallote):

what do you think is the second thing we dont know?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what x is?

OpenStudy (lgbasallote):

nope...that's not really important

OpenStudy (lgbasallote):

look at the product rule

OpenStudy (anonymous):

oh

OpenStudy (lgbasallote):

what's the second thing you dont know in the product rule?

OpenStudy (anonymous):

d/dx?

OpenStudy (lgbasallote):

d/dx of what?

OpenStudy (anonymous):

x^2+1

OpenStudy (lgbasallote):

right. so what's \[\frac{d}{dx} (x^2 + 1)\]

OpenStudy (anonymous):

2x

OpenStudy (lgbasallote):

right so now substitute

OpenStudy (lgbasallote):

substitute the value of \[\frac{d}{dx} \sin \sqrt x\] and the value of \[\frac{d}{dx} (x^2 + 1)\] into \[(x^2 + 1) \frac{d}{dx} (\sin \sqrt x) + \sin \sqrt x \frac{d}{dx} (x^2 + 1)\]

OpenStudy (anonymous):

2x(1/2 cos ^-1/2)+(1/2 cos ^-1/2)(2x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont think that is right

OpenStudy (anonymous):

would it be (x^2+1)(1/2 cos ^-1/2)+(sin sqrtx)2x

OpenStudy (lgbasallote):

what did you do?

OpenStudy (anonymous):

substituted

OpenStudy (lgbasallote):

where are the x

OpenStudy (anonymous):

huh?

OpenStudy (lgbasallote):

(x^2+1)(1/2 cos ^-1/2)+(sin sqrtx)2x where's the x on that cos thingy

OpenStudy (anonymous):

would it be (x^2+1)(1/2 cosx ^-1/2)+(sin sqrtx)2x

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

can you draw it?

OpenStudy (anonymous):

draw what?

OpenStudy (lgbasallote):

your answer

OpenStudy (anonymous):

\[x ^{2}+1(1/2\cos ^{-1/2})+(\sin \sqrt{x})(2x)\]

OpenStudy (lgbasallote):

no...not the equation...the draw tbutton...draw your answer

OpenStudy (anonymous):

isnt that my answer?

OpenStudy (lgbasallote):

because there's no such thing as \[\cos^{-1/2}\]

OpenStudy (lgbasallote):

im looking for x...idk where you placed it

OpenStudy (lgbasallote):

look at what i said \[\frac{d}{dx} \sin \sqrt x\] again

OpenStudy (anonymous):

cosx^-1/2

OpenStudy (lgbasallote):

what else? that's incomplete

OpenStudy (lgbasallote):

look at the post where i said what d/dx sin sqrt x is

OpenStudy (anonymous):

do you mean the 1/2 infront?

OpenStudy (lgbasallote):

what did i say d/dx sin sqrt x was?

OpenStudy (anonymous):

cos√x /2√x

OpenStudy (lgbasallote):

can you see the difference with what you wer saying a while ago?

OpenStudy (anonymous):

yupp

OpenStudy (lgbasallote):

so what's the complete correct answer now?

OpenStudy (anonymous):

\[x ^{2}+1(\cos \sqrt{x}/2\sqrt{x})+\sin \sqrt{x}(2x)\]

OpenStudy (lgbasallote):

right \[\large \frac{(x^2+1)(\cos \sqrt x)}{2\sqrt x} + 2x\sin \sqrt x\]